two years ago father was 3 times as old as his son and 2 year hence,twice his age will be equal to 5 times that of his son find present age ?
Answer by Maths68(1474)

Let
Present age of Father = f
Present age of Son = s

Two years ago
f-2=s-2
Father was 3 times as old as his son
f-2=3(s-2)
f-2=3s-6
f=3s-6+2
f=3s-4............(1)
and 2 year hence,
f+2=s+2
then twice his age will be equal to 5 times that of his son
2(f+2)=5(s+2)
2f+4=5s+10
2f=5s+10-4
2f=5s+6.............(2)
Put the value of f in (2) from (1)
2(3s-4)=5s+6
6s-8=5s+6
6s-5s=6+8
s=14
Put the value of s in (1)
f=3s-4............(1)
f=3(14)-4
f=42-4
f=38
Present age of Father = f = 38 years old
Present age of Son = s = 14 years old